The first term of a geometric progression is 5 and the fifth term is 1280, the common ratio is 4
Find the expression of the nth term?
Can someone explain this step by step please

Answer by stanbon(75887)


The first term of a geometric progression is 5 and the fifth term is 1280, the common ratio is 4
Find the expression of the nth term?
--------
a(1) = 5
r = 4
----
a(5) = 5*4^x = 1280
4^x = 256
x = 4
So the exponent of the common ratio is one-less than the number of the term.
------
Ans: a(n) = 5*4^(n-1)
------------------
Cheers,
Stan H.
--------------